OpenStudy (anonymous):

What is -5/8 divided by 9/10? A. -25/36 B. - 9/16 C. -45/80 D. -1/36

OpenStudy (anonymous):

I think it's C

OpenStudy (nikita2):

(-5/8)/(9/10) = - ((5*10)/(8*9)) = -25/36

OpenStudy (anonymous):

Thank you